# Double Higgs Production via Vector Boson Fusion in SMEFT

### v1

@@ A.3.2 $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$, $X\varphi^{4}D^{2}$ operator classes
±±$\mathcal{M}^{WW,\,D8}_{\pm\pm}=$ $-\mathopen{}\mathclose{{}\left(\frac{1}{4}C_{W^{2}\varphi^{2}D^{2}}^{(1)}+C_{W^{2}\varphi^{2}D^{2}}^{(2)}}\right)\mathopen{}\mathclose{{}\left(\frac{s^{2}}{\Lambda^{4}}}\right)$
@@
±$\mathcal{M}^{WW,\,D8}_{\pm\mp}=$ $-\frac{1}{8}\sin^{2}\theta C_{W^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{}\left(\frac{s^{2}}{\Lambda^{4}}}\right)+2^{3/4}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{}\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}\right)\mathopen{}\mathclose{{}\left(\frac{M_{W}^{2}}{\Lambda^{2}}}\right)\;,$
@@ Appendix B $ZZ\rightarrow HH$ helicity amplitude
±±$\mathcal{M}^{ZZ,\,SM}_{\pm\pm}$ $=\mathcal{O}\mathopen{}\mathclose{{}\left(\frac{1}{s}}\right)\;,$
±$\mathcal{M}^{ZZ,\,SM}_{\pm\mp}$ $=2\sqrt{2}G_{F}M_{Z}^{2}\;,$
$\mathcal{M}^{ZZ,\,SM}_{0\pm}$ =±$=\mathcal{M}^{ZZ,\,SM}_{\pm 0}=\mathcal{O}\mathopen{}\mathclose{{}\left(\frac{1}{\sqrt{s}}}\right)\;.$
@@ B.2 Dimension-6 SMEFT
[…]
@@
±±$\mathcal{M}^{ZZ,\,D6}_{\pm\pm}=$ $-2(2C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{}\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}\right)$
@@
±$\mathcal{M}^{ZZ,\,D6}_{\pm\mp}=$ $+(4C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{}\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}\right)$
@@ B.3.1 $\varphi^{8}$, $\varphi^{6}D^{2}$, $\varphi^{4}D^{4}$ operator classes
±±$\mathcal{M}^{ZZ,\,D8}_{\pm\pm}=$ $-2\sqrt{2}\mathopen{}\mathclose{{}\left(2C_{\varphi^{6}\Box}+C_{\varphi^{6}D^{2}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}$
@@
±$\mathcal{M}^{ZZ,\,D8}_{\pm\mp}=$ $+\sqrt{2}\mathopen{}\mathclose{{}\left(2C_{\varphi^{6}\Box}+\frac{3}{4}C_{\varphi^{6}D^{2}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}+\frac{1}{2}C_{\varphi^{4}D^{4}}^{(2)}\sin^{2}\theta\mathopen{}\mathclose{{}\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}\right)\mathopen{}\mathclose{{}\left(\frac{s}{\Lambda^{2}}}\right)\;,$
@@ B.3.2 $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$, $X\varphi^{4}D^{2}$ operator classes
±±$\mathcal{M}^{ZZ,\,D8}_{\pm\pm}=$ $+3\sqrt{2}C_{B^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{}\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\right)\mathopen{}\mathclose{{}\left(\frac{1}{G_{F}\Lambda^{2}}}\right)\mathopen{}\mathclose{{}\left(\frac{s}{\Lambda^{2}}}\right)+3\sqrt{2}C_{W^{2}\varphi^{4}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\mathopen{}\mathclose{{}\left(\frac{1}{G_{F}\Lambda^{2}}}\right)\mathopen{}\mathclose{{}\left(\frac{s}{\Lambda^{2}}}\right)$
@@
±$\mathcal{M}^{ZZ,\,D8}_{\pm\mp}=$ $-\frac{1}{8}C_{B^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{}\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}\right)\sin^{2}\theta\mathopen{}\mathclose{{}\left(\frac{s^{2}}{\Lambda^{4}}}\right)-\frac{1}{8}C_{W^{2}\varphi^{2}D^{2}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\sin^{2}\theta\mathopen{}\mathclose{{}\left(\frac{s^{2}}{\Lambda^{4}}}\right)$

### v2

@@ A.3.2 $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$, $X\varphi^{4}D^{2}$ operator classes
$\mathcal{M}^{WW,\,D8}_{\pm\pm}=$ $-\mathopen{}\mathclose{{\left(\frac{1}{4}C_{W^{2}\varphi^{2}D^{2}}^{(1)}+C_{W^{2}\varphi^{2}D^{2}}^{(2)}}}\right)\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)$
@@
$\mathcal{M}^{WW,\,D8}_{\pm\mp}=$ $-\frac{1}{8}\sin^{2}\theta C_{W^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)+2^{3/4}C_{W\varphi^{4}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(\frac{M_{W}}{\sqrt{G_{F}}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{M_{W}^{2}}{\Lambda^{2}}}}\right)\;,$
@@ Appendix B $ZZ\rightarrow HH$ helicity amplitude
$\mathcal{M}^{ZZ,\,SM}_{\pm\pm}$ $=\mathcal{O}\mathopen{}\mathclose{{\left(\frac{1}{s}}}\right)\;,$
$\mathcal{M}^{ZZ,\,SM}_{\pm\mp}$ $=2\sqrt{2}G_{F}M_{Z}^{2}\;,$
$\mathcal{M}^{ZZ,\,SM}_{0\pm}$ $=\mathcal{M}^{ZZ,\,SM}_{\pm 0}=\mathcal{O}\mathopen{}\mathclose{{\left(\frac{1}{\sqrt{s}}}}\right)\;.$
@@ B.2 Dimension-6 SMEFT
$\mathcal{M}^{ZZ,\,D6}_{00}=$ $-3\sqrt{2}C_{\varphi}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)+(2C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\hskip 190.63338pt\textbf{dim-6}$
@@
$\mathcal{M}^{ZZ,\,D6}_{\pm\pm}=$ $-2(2C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)$
@@
$\mathcal{M}^{ZZ,\,D6}_{\pm\mp}=$ $+(4C_{\varphi\Box}+C_{\varphi D})\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)$
@@ B.3.1 $\varphi^{8}$, $\varphi^{6}D^{2}$, $\varphi^{4}D^{4}$ operator classes
$\mathcal{M}^{ZZ,\,D8}_{\pm\pm}=$ $-2\sqrt{2}\mathopen{}\mathclose{{\left(2C_{\varphi^{6}\Box}+C_{\varphi^{6}D^{2}}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}$
@@
$\mathcal{M}^{ZZ,\,D8}_{\pm\mp}=$ $+\sqrt{2}\mathopen{}\mathclose{{\left(2C_{\varphi^{6}\Box}+\frac{3}{4}C_{\varphi^{6}D^{2}}}}\right)\frac{M_{Z}^{2}}{\Lambda^{2}}\frac{1}{G_{F}\Lambda^{2}}+\frac{1}{2}C_{\varphi^{4}D^{4}}^{(2)}\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{M_{Z}^{2}}{\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)\;,$
@@ B.3.2 $X^{2}\varphi^{4}$, $X^{2}\varphi^{2}D^{2}$, $X\varphi^{4}D^{2}$ operator classes
$\mathcal{M}^{ZZ,\,D8}_{\pm\pm}=$ $+3\sqrt{2}C_{B^{2}\varphi^{4}}^{(1)}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)+3\sqrt{2}C_{W^{2}\varphi^{4}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\mathopen{}\mathclose{{\left(\frac{1}{G_{F}\Lambda^{2}}}}\right)\mathopen{}\mathclose{{\left(\frac{s}{\Lambda^{2}}}}\right)$
@@
$\mathcal{M}^{ZZ,\,D8}_{\pm\mp}=$ $-\frac{1}{8}C_{B^{2}\varphi^{2}D^{2}}^{(1)}\mathopen{}\mathclose{{\left(1-\frac{M_{W}^{2}}{M_{Z}^{2}}}}\right)\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)-\frac{1}{8}C_{W^{2}\varphi^{2}D^{2}}^{(1)}\frac{M_{W}^{2}}{M_{Z}^{2}}\sin^{2}\theta\mathopen{}\mathclose{{\left(\frac{s^{2}}{\Lambda^{4}}}}\right)$